Differences between revisions 18 and 19
Revision 18 as of 2022-08-05 18:29:38
Size: 2844
Revision 19 as of 2022-08-05 18:40:59
Size: 2854
Deletions are marked like this. Additions are marked like this.
Line 30: Line 30:
 * TMS: Pressure difference in EX.  * TMS: Replacement of pressure gauge in EX.


Zoom Meeting

CRY+VIS Interface Meeting

Date: 2022/8/5 14:00-14:20

Participants: Takahashi, Miyoki, Ushiba, Tanaka, Sato, Kimura, Ikeda, Aso, Akutsu

Post O3 Upgrade portal


  • Type-A: Investigation of ADC/BIO noises.
  • Type-B: Damping of SR2/3 and BS. Replacement of main card for LVDT driver in SR2/3. Cable check. Preparation for review in the next week.
  • Type-Bp: Height check. Review on 8/5.

  • Type-C:


  • Replacement of cables: EXT and IY/EX remain.
  • TM height setting: Type-Bp
  • Oplev: Investigation of the glitches.
  • ACC:
  • Commissioning in air:
  • Commissioning in vacuum: ALS DARM was locked.
  • Acceptance check: Dropbox paper

  • New vacuum pumps: Installation of ion pumps in IX/IY on 8/4-5. Side hatch in SRM on 8/22.
  • Heater control: Pending.
  • TMS: Replacement of pressure gauge in EX.
  • Cooling: Started in IY.


VIS Meeting

Date: 2022/8/5 14:30-14:45

Participants: Takahashi, Sato, Ushiba, Aso, Fabian, Tanaka, Hirata

Summary Report



  • Hirata-san and others replaced the master boards in the LVDT drivers in SR2 and SR3. In both cases there were reductions of the noise, especially at 10 Hz.
  • Hirata-san and others also checked again the integrity of the cables in SR suspensions. There were some concerns that we checked today:
    • In SR2, pin 8 of the IM yaw picomotor has a connection to the ground pin inside the chamber. This is new, but not harmful because such a pin is connected to the ground in the driver.
    • The ground pin of the break-out cable is not so stable, so we were measuring higher values than before. Instead, we measured resistances by touching the vacuum chamber instead.
    • We also found that some resistances between some pins in the SR3 F1 fishing rod, increased by 20 Ohm or so. We don't know the reason, but it's still fine because we are using the stand alone driver.


  • Height check.
  • Review.





  • Type-A:
  • Type-B: Health check. Preparation for review.
  • Type-Bp:
  • Type-C:



Travel Plans

(Date in Kamioka)

The week from 8/8 (Obon 8/11-16)

  • Takahashi: 8/9-10
  • Hirata: 8/8-10

The week from 8/15

  • Takahashi: 8/14-17 (safety shift)
  • Hirata: 8/16-19
  • Aso: 8/16-19

Next meeting

  • 2022/8/19(FRI)

KAGRA/Subgroups/VIS/VISMinutes20220805 (last edited 2022-08-05 18:40:59 by RyutaroTakahashi)